OpenStudy (anonymous):

How do I Solve the equation 4/3t=5t/7

OpenStudy (anonymous):

\[\frac{ 4 }{ 3t }=\frac{ 5t }{ 7 }\] \[4=\frac{ 15t ^{2} }{ 7 }\] We multiplied the 3t by the 5t \[28=15t ^{2}\] \[t \approx 1.36 \] I think thats how its solved

OpenStudy (anonymous):

Almost, but t is not 1.36, you may need to resolve the t after 28 = 5t^2